alones (alones):

An object is dropped off a building that is 144 feet tall. After how many seconds does the object hit the ground?

alones (alones):

(s = 16t^2)

OpenStudy (kendricklamar2014):

Use the drt formula: \[distance = rate \times time \]

OpenStudy (the_puppet1):

I think about 6 seconds until it hits the ground

OpenStudy (the_puppet1):

When it is just dropped just count the sec so the answer is 6 seconds

OpenStudy (shadowlegendx):

This is ignoring other factors right? Like air resistance

OpenStudy (kendricklamar2014):

Now just substitute: \[144 = 16t^2 * X \]

alones (alones):

Okay i got \[\frac{ 9 }{ t^2 }\]

OpenStudy (kendricklamar2014):

\[s = 16t^2\] \[144 = 16t^2\] \[9 = t^2\] What do you get?

OpenStudy (mjdennis):

@AloneS , do back to your first equation. s is distance, in this case height. So solve 144 = 16*t^2 for t (and throw away the negative answer, since t^2=something has two answers...)

OpenStudy (kendricklamar2014):

@AloneS To solve this problem, all you need to do is find what number squared is equal to 9. Do you understand what I mean by that?

alones (alones):

No sorry i'm confused now

OpenStudy (kendricklamar2014):

Okay, so we need to substitute t^2. What number can we substitute so that x(the number(^2 is equal to 9.

alones (alones):

Uhh 3?

OpenStudy (kendricklamar2014):

Yes, that is correct =)

alones (alones):

YAAAASSS!!!!!FINALLYYYY!!!THANSK KENDRIC<3

OpenStudy (kendricklamar2014):

Your welcome :)

alones (alones):

THNAKS<3